Simple-RegressionTech: Problem 1 Previous Problem Problem List Next Problem (1 point) Measuring Temperature. The two most commonly used scales for measuring temperature are the Fahrenheit and Celsius scales. If you let y denote the Fahrenheit temperature and x denote Celsius temperature, you can express the relationship between those two scales with the linear equation y # 32 + 1.8x. a. Determine the y-intercept bo and the slope b .Note that a and b of the lecture notes are b-0 and b-1 here. b. Find the Fahrenheit temperature corresponding to the Celsius temperature -16 degrees c. Find the Fahrenheit temperature corresponding to the Celsius temperature O degrees d. Find the Fahrenheit temperature corresponding to the Celsius temperature 14. degrees e. Find the Fahrenheit temperature corresponding to the Celsius temperature 100 degrees f. Graph the linear equation y 32+1.8x, using the four points found in (b), (c). (d) and (e). g. Apply the graph obtained in part (0) to estimate visually the Fahrenheit temperature corresponding to a Celsius temperature of 23. Then calculate that temperature exactly by using the linear equation y - 32 +1.8x. The exact Fahrenheit temperature corresponding to the Celsius temperature of 23° is degrees
Answer #1

The linear regression equation is y = 32 + 1.8*x

a)

b0 = 32

b1 = 1.8

b)

x = - 16

y = 32 + 1.8*(-16) = 3.2

c)

x = 0

y = 32 + 1.8*0 = 32

d)

x = 14

y = 32 + 1.8*14 = 57.2

e)

x = 100

y = 32 + 1.8*100 = 212